# A Comprehensive Study on Large Language Models for Mutation Testing

### v3

@@ 6.2. LLMs for Code
Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023). Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes. LLMs are driven by instructions in natural languages, i.e., prompts, which guide the models to perform target tasks and are crucial to performance. The design of prompts plays a crucial role in obtaining accurate and useful responses from these models.
@@ 7. Conclusion
In this paper, we systematically investigate the performance of existing mutation generation appraoches across multiple dimensions, including their effectiveness, validity, and efficiency. Our study conducts extensive and detailed comparative experiments, covering both widely studied rule-based mutation generation approaches and LLM-based methods. Our evaluation results highlight that LLMs significantly outperform traditional approaches in generating diverse mutations that closely mimic real bugs. However, the results also expose key limitations of LLMs, particularly their tendency to generate a high proportion of non-compilable mutations. Therefore, we analyze the error types of the LLM-generated non-compilable mutations and identify member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations. Based on our findings, we emphasize the need for further research to enhance LLM-based mutation testing, addressing its current shortcomings while leveraging its strengths.

### v4

@@ 6.2. LLMs for Code
Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023b). Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes. LLMs are driven by instructions in natural languages, i.e., prompts, which guide the models to perform target tasks and are crucial to performance. The design of prompts plays a crucial role in obtaining accurate and useful responses from these models.
@@ 7. Conclusion
In this paper, we systematically investigate the performance of existing mutation generation appraoches across multiple dimensions, including their effectiveness, validity, and efficiency. Our study conducts extensive and detailed comparative experiments, covering both widely studied rule-based mutation generation approaches and LLM-based methods. Particularly, we evaluate not existing LLM-based approaches but also our new prompt for mutation generation. Our evaluation results highlight that LLMs significantly outperform traditional approaches in generating diverse mutations that closely mimic real bugs. However, the results also expose key limitations of LLMs, particularly their tendency to generate a high proportion of non-compilable mutations. Therefore, we analyze the error types of the LLM-generated non-compilable mutations and identify that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations. Based on our findings, we emphasize the need for further research to enhance LLM-based mutation testing, addressing its current shortcomings while leveraging its strengths.